OpenStudy (anonymous):

Three resistors are connected in parallel in a circuit with a 9-volt battery. The resistors are 4, 6 and 12 ohms. Calculate the current in the circuit.

OpenStudy (anonymous):

@supercrazy92 can u help?

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you know how to calculate the resistance in parallel combination ?

OpenStudy (anonymous):

well i added them and i got 22 and i divided by 9 but i dont know if i got it right

OpenStudy (anonymous):

the answer choices are these A 18 amps B. 9 amps C. 4.5 amps D. 2.25 amps

OpenStudy (anonymous):

For parallel combination \[\frac{ 1 }{ R } = \frac{ 1 }{ R_1 } + \frac{ 1 }{ R_2 } + \frac{ 1 }{ R_1 }\] Solve this , you'll get the Total resistance after that , use Ohm's law again to find I (current).

OpenStudy (anonymous):

so the total resistance is 22

OpenStudy (anonymous):

i got .40909

OpenStudy (anonymous):

ohh nvm i think i got it it will be 18

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

No, it should be 2ohm

OpenStudy (anonymous):

but its not on the answer choices :/

OpenStudy (anonymous):

.409 that must be 1/R for R you should divide this (1/.409)

OpenStudy (anonymous):

well the only 2 in the answer choices is 2.25 and that says 2.44

OpenStudy (anonymous):

I got 1/R =0.499 and R =2 ohm I told you it is not the answer it is the totalt resistance of the circuit. now we have to use Ohm's law for total current the circuit.

OpenStudy (anonymous):

Did you get that ?

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it will be 4.5 ohms

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

V= IR I = V/R

OpenStudy (anonymous):

amps i mean

OpenStudy (anonymous):

yes, exactly ;)

OpenStudy (anonymous):

i got it thanks!